Simplify completely: (e4f/g) divided by (ef3/g3)
Answer by jayanthihemkos@yahoo.com(87)

Simplify completely: (e4f/g) divided by (ef3/g3)
==> (e^4f/g)/(ef^3/g^3)

==> (e^4fg^-1/ef^3g^-3) [a^-m = 1/a^m]
==> e^4-1f^1-3g^-1+3 [ a^m/a^n = a^m-n)
==> e^3f^-2g^2

==> e^3g^2/f^2 [a^-m = 1/a^m]